OpenStudy (anonymous):

(n+1)!/(n-1)!=20 n!=?

OpenStudy (a_clan):

the expression (n+1)!/(n-1)! should be part of an equality

OpenStudy (a_clan):

(n+1)!/(n-1)! = RHS

OpenStudy (anonymous):

RHS?

OpenStudy (a_clan):

right hand side (LHS) x = y (RHS)

OpenStudy (anonymous):

try n^2+n put values

OpenStudy (anonymous):

(n+1)!/(n-1)!=20 n!=?

OpenStudy (anonymous):

any1?

OpenStudy (anonymous):

like I said, try n^2+n

OpenStudy (a_clan):

The trivial method is to find 'n' and then solve n! , as panlac01 said. Method 1: (n+1)!/(n-1)!=20 (n+1)n. (n-1)! / (n-1)! = 20 (n+1)n = 20 or n^2 + n = 20

OpenStudy (anonymous):

solve for N for the answer?

OpenStudy (a_clan):

OR Method 2: (n+1)!/(n-1)! = (n+1) n! / (n-1)! .....................eqn (A) || 20 = 5*4 = 5 * 4 *(3!) / (3!) [Multiply and divide by 3!] = 5 * 4! / 3! [Now this is the same form as the RHS of eqn (A)] You can compare n! = 4! =24 (SOLUTION}

OpenStudy (anonymous):

thank u. the "!" confuesd me. thanx for the help, bro!

OpenStudy (a_clan):

np